OpenStudy (anonymous):

Find an equation of he tangent line to the curve at the given point y= sinx+sin^2 x , (0,0)

OpenStudy (anonymous):

y' = cos(x) + cos(x)sin(x) + sin(x)cos(x) y' = cos(x) + 2cos(x)sin(x) At (0,0), y' = cos(0) + 2cos(0)sin(0) = 1 There's your slope. Now for the equation, y + 0 = 1(x+0) y=x

OpenStudy (anonymous):

Well, to find the slope of an equation at any point in the function, we first have to find the derivative of the function. The derivative of y = sin(x) + sin^2(x) is y' = cos(x) + 2cos(x)sin(x). To find the slope at a specific point, you just plug in the x-value of the point into the derivative. I plugged in zero, and got the slope, then used point-slope formula to get the equation of the line. What exactly don't you understand?

OpenStudy (anonymous):

its finding the derivatives taht i find diffivcult to do

OpenStudy (anonymous):

Okay. We have y = sin(x) + sin^2(x) In order to find the derivative of this function, we must know that d/dx [sin(x)] = cos (x). Then, we must know that d/dx [f(x) + g(x)] = d/dx [f(x)] + d/dx [g(x)]. And also the product rule: d/dx [f(x) * g(x)] = f'(x)g(x) + f(x)g'(x) So, y' = d/dx [sin(x) + sin^2(x)] y' = d/dx [sin(x)] + d/dx [sin^2(x)] The derivative of sin(x) is cos(x). y' = cos(x) + d/dx [sin^2(x)] Now, let's deal with sin^2(x). This can be expanded to sin(x) * sin(x). Product rule. d/dx [ sin(x)*sin(x) ] = d/dx [sin(x)]sin(x) + d/dx [sin(x)]sin(x) = cos(x)sin(x) + cos(x)sin(x) = 2cos(x)sin(x).

OpenStudy (anonymous):

thankls foir breaking ut down....it makes it easier b.c wen i try to undertans math n steps are skipped it confuses me...srrry im just slow in math